ACTION 18 ttdf1. Accrued interest Payable Compute the interest accrued on each of the foilowing notes payable owed bay Gsloway Galloway, Inc, on December 31 Lender Date of Note Principal Interest Rate (%) Term Barton12/4$10000 Lawson 1213 14000 Rlley 12/1915.000 12 150 days 9 90 days 11 60 days Lender Accrued Interest Ridey de OEST?ON 19 Not yet answered Marked aldan if a company issues 10,000 shares of the following is the correct balance sheet entry? 54 par value common stock at a market price of $60 per share, which of land ho Select one: A Increase cash by $600,000 and Increase pald-in capital by $600.000 ?B. increase cash by $600,000 and increase retained earnings by S600.000 al i O D. Increase common stock and cash by $40,000 QUESION 20 Notyeranswered ed u of 2.0 On September 1, 2016, Caroline Companys balance sheet indicates there are 1,200,000 shares of $60 par value common shares in the Common Stock account and $59,000,000 in the Additional Paid-in Capital account. There are 2,000,000 shares authorized. On September 2, Caroline splits its stock 2 for 1 What is the dollar balance of Carolines cammon stack account Immediatcly after thc stock spli Select one: A$36,000,00 B$72,000,000 o C.$ 18,000,000 o D. $144,000,000 o trom iniluencing the calculation of the ratio,
Answer #1

18). Calculation of Accrued Interest on Dec 31
Barton = $10000* 12% = $1200/365 *27 days = $88.77
Lawson = $14000 * 9% = $1260/365*18 days = $62.14
Riley = $15000 * 11% = $1650/365*12 days = $54.25

19). Entry for issuing common stock of $4 par value at $60 market value.

Account Titles Debit Credit
Cash $600000
Common Stock at Par Value $40000
Paid in Capital in excess of par value $560000

Hence answer is A.

20). Balance of Common stock account = 1,200,000 shares * $60 par value = $72,000,000
This balance remains the same if there is stock split because there will be no change in the account balance. Decrease in par value of stock is offset by increase in no. of shares. Hence the account balance remains same.